OpenStudy (anonymous):

Solve the equation. 1/3 log2(x + 6) = log8(3x)

OpenStudy (turingtest):

\[{1\over3} \log_2(x + 6) = \log_8(3x)\]right?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

\[{1\over3}\log_2(x+6)={\log_23x \over \log_28}={\log_23x \over3}\]can you take it from here?

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

change of base is good. also note that if you exponentiate with base 8 you get \[8^{\log_8(3x)}=8^{\frac{1}{3}\log_2(x+6)}\] \[3x=2^{\log_2(x+6)}\] \[3x=x+6\]

OpenStudy (turingtest):

So you do :)

OpenStudy (turingtest):

Ah sat's a mod now too? they must have made a round...

OpenStudy (anonymous):

i think it just comes with the medals or something. never volunteered.